If a perpetuity that pays 10 once a year, what is its Macaulay duration? Assume a 5% yield:

A. 22

B. 200

C. 210

D. 21

Please explain your answer.
Answer #1

Annual Payment = $10

Annual yield = i = 5% or 0.05

Macaulay Duration = (1+i)/i = (1+0.05)/0.05 = 21